# label1]organization=Consejo Superior de Investigaciones Científicas, Centre for Automation and Robotics-CAR (CSIC-UPM), addressline=Ctra. Campo Real km. 0,200, postcode=28500, state=Arganda, country=Spain

### v1

@@ Abstract
The effective integration of generative artificial intelligence in education is a fundamental aspect to prepare future generations. This study proposes an accelerated learning methodology in artificial intelligence, focused on its generative capacity, as a way to achieve this goal. It recognizes the challenge of getting teachers to engage with new technologies and adapt their methods in all subjects, not just those related to AI. This methodology not only promotes interest in science, technology, engineering and mathematics, but also facilitates student understanding of the ethical uses and risks associated with AI. Students’ perceptions of generative AI are examined, addressing their emotions towards its evolution, evaluation of its ethical implications, and everyday use of AI tools. In addition, AI applications commonly used by students and their integration into other disciplines are investigated. The study aims to provide educators with a deeper understanding of students’ perceptions of AI and its relevance in society and in their future career paths.
@@ 1 Introduction
2 Methodological framework and information retrieval
It is presented a detailed proposal to teach artificial intelligence to middle and high school students, focusing on its ethical uses and practical applications in their daily lives. This educational program is structured in two main segments to ensure a deep and thoughtful understanding of the subject. In the first part, the basics of AI are addressed, starting with the definition of human intelligence. This section is designed to be formative and dynamic, lasting approximately one hour. During this initial session, strategic questions are posed to students to encourage reflection on the concept of AI and its roots in human intelligence. They are asked what they understand by intelligence. Student responses included concepts such as "ability to solve problems, analyze, reason, innovate, knowledge, logic and critical reasoning, think”. After capturing the students’ attention and initiating an active dialogue with them, exploration of the concept of artificial intelligence commenced. Despite being a term widely used in today’s society, it is observed that it has not been analyzed in depth in its close context. The students’ answers to the question of what they understood by artificial intelligence included descriptions such as “problem-solving machine, imitating human intelligence, improvement of human intelligence, the ability of non-living beings to solve problems, non-human being with the intention of copying us”. It is remarkable how many students perceive artificial intelligence as a tool intended to enhance human capability.
@@ 3 Experiment and results
Figure 4: Videos produced by students’ free prompts. a) A man in egypt with pyraminds in the background, a man in the ancient greece, a roman emperor during the fall of roman empire, a native american being colonized, a french republican during french revolution. b)Primitive man ans women in a cavern, man and women of the roman empire, man and women archer in a castle, mustketeer and a woman musketeer in the medium age, man and women in the wall street of new york. c) a prehistorical landscape with a bunch of dinosaurs trending on artstation, four pyramids in the dessert with a sunny sky, the roman coloseum with a gradiator fighing against a lion, an historical castle with with a rtunch under a storm, a city of Dubai in a sunny day with their skyscrapers. d) a beautiful city with cubism style, a beautiful city with Van Gogh style, a beautiful city with Da Vinci style, a beautiful city with Velazquez style, a beautiful city with Picasso style. e) Fractals, vivid colors. f) a beautiful boy dressed like 1920s style, a beautiful boy dressed like 1940s style, a beautiful boy dressed like 1960s style, a beautiful boy dressed like a rock style, a beautiful boy dressed with sport clothes. g) Man and a woman dressed like in the 1920s, Man and a woman dressed like in the 1940s, Man and a woman dressed like in the 1960s, Man and a woman dressed like in the 1980s, Man and a woman dressed like in the 2000s.
@@ 4 Research questions discussion
Despite the educational potential of AI, 90% of students indicated that their teachers do not encourage them to use AI technologies as support tools in their classes. This points to a significant gap between students’ interest and willingness to use AI and support or promotion by educators. The lack of incentive on the part of teachers suggests an opportunity for improvement in the integration of AI in teaching, highlighting the need to educate and train teachers in these emerging technologies to maximize their potential in the classroom.
@@ 5 Conclusions
6 Acknowledgement
This activity is part of the 4ºESO+EMPRESA Program organized by the Subdirección General de Centros de Educación Secundaria de la Comunidad de Madrid, in collaboration with the CSIC Delegation in Madrid and organized and developed by the Artificial Perception Group of the Centro de Automática y Robótica. Héctor Montes is grateful to the National Research System of Panama (SNI-SENACYT) for the support provided to its members.

### v2

@@ Abstract
The effective integration of generative artificial intelligence in education is a fundamental aspect to prepare future generations. The objective of this study is to analyze from a quantitative and qualitative point of view the perception of controlled student-IA interaction within the classroom. This analysis includes assessing the ethical implications and everyday use of AI tools, as well as understanding whether AI tools encourage students to pursue STEM careers. Several points for improvement in education are found, such as the challenge of getting teachers to engage with new technologies and adapt their methods in all subjects, not just those related to technologies.
@@ 1 Introduction
2 AI - student interaction and information retrieval
A controlled framework is presented in which secondary school students are exposed to the use of generative AI for subsequent analysis, focusing on its ethical uses and practical applications in their daily lives. This educational program is structured in two main segments to ensure a deep and thoughtful understanding of the subject. In the first part, the basics of AI are addressed, starting with the definition of human intelligence. This section is designed to be formative and dynamic, lasting approximately one hour. During this initial session, strategic questions are posed to students to encourage reflection on the concept of AI and its roots in human intelligence. They are asked what they understand by intelligence. Student responses included concepts such as "ability to solve problems, analyze, reason, innovate, knowledge, logic and critical reasoning, think”. After capturing the students’ attention and initiating an active dialogue with them, exploration of the concept of artificial intelligence commenced. Despite being a term widely used in today’s society, it is observed that it has not been analyzed in depth in its close context. The students’ answers to the question of what they understood by artificial intelligence included descriptions such as “problem-solving machine, imitating human intelligence, improvement of human intelligence, the ability of non-living beings to solve problems, non-human being with the intention of copying us”. It is remarkable how many students perceive artificial intelligence as a tool intended to enhance human capability.
@@ 3 Experiment and results
Figure 4: Videos produced by students’ free prompts. a) A man in egypt with pyraminds in the background, a man in the ancient greece, a roman emperor during the fall of roman empire, a native american being colonized, a french republican during french revolution. b)Primitive man ans women in a cavern, man and women of the roman empire, man and women archer in a castle, mustketeer and a woman musketeer in the medium age, man and women in the wall street of new york. c) a prehistorical landscape with a bunch of dinosaurs trending on artstation, four pyramids in the dessert with a sunny sky, the roman coloseum with a gradiator fighing against a lion, an historical castle with with a trench under a storm, a city of Dubai in a sunny day with their skyscrapers. d) a beautiful city with cubism style, a beautiful city with Van Gogh style, a beautiful city with Da Vinci style, a beautiful city with Velazquez style, a beautiful city with Picasso style. e) Fractals, vivid colors. f) a beautiful boy dressed like 1920s style, a beautiful boy dressed like 1940s style, a beautiful boy dressed like 1960s style, a beautiful boy dressed like a rock style, a beautiful boy dressed with sport clothes. g) Man and a woman dressed like in the 1920s, Man and a woman dressed like in the 1940s, Man and a woman dressed like in the 1960s, Man and a woman dressed like in the 1980s, Man and a woman dressed like in the 2000s.
@@ 4 Research questions discussion
Despite the educational potential of AI, 90% of students indicated that their teachers do not encourage them to use AI technologies as support tools in their classes. It points to a significant gap between students’ interest and willingness to use AI and support or promotion by educators. The lack of incentive on the part of teachers suggests an opportunity for improvement in the integration of AI in teaching, highlighting the need to educate and train teachers in these emerging technologies to maximize their potential in the classroom.